# MInference 1.0: Accelerating Pre-filling for Long-Context LLMs via Dynamic Sparse Attention

### v1

@@ 1 Introduction
Large language models (LLMs) have entered the era of long-context processing, with some of them supporting context windows ranging from 128K to 10M tokens [24, 67, 49, 84, 2, 12]. These extended context windows enable LLMs to unlock a multitude of complex real-world applications, such as repository-level code understanding [7, 34, 58], long-document question-answering [9, 51], extreme-label in-context learning [51], and long-horizon agent tasks [79].
@@ 2.2 Attention Sparsity Exhibits Patterns
Vertical-Slash (VS) pattern The attention weights are concentrated on specific tokens (vertical lines) and tokens at fixed intervals (slash lines). The positions of vertical and slash lines in this pattern dynamically change with the context content and exhibit a certain sparsity, making them difficult to be encompassed by local windows and A-shape patterns.
@@ 4 Experiments
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Number Retr.KV Avg.
@@
As shown in Table 2, MInference achieves the best overall performance compared to baselines on InfiniteBench. Remarkably, MInference matches or even slightly surpasses the performance of the original full attention baseline on some tasks, despite the significant acceleration it provided. From the perspective of different tasks, our method not only performs well in natural language tasks such as summarization, QA, and code, but also maintains the original model’s performance on retrieval-related tasks. Baseline methods such as StreamingLLM, on the contrary, struggle with these retrieval tasks. Additionally, on tasks such as dialogue QA, using local attention mechanisms can better handle these tasks, while our performance is closer to the original results, indicating that our method is not solely based on local windows. Extending the local windows’ intervals in StreamingLLM, i.e., w/ dilated and w/ strided, hardly affects the model’s performance.
@@
Comparing Fig. 1a to 6, our method effectively retains the ability to process information at different positions across various context windows, ranging from 1k to 1M tokens. In contrast, methods like StreamingLLM, while effective in reducing latency, experience a rapid decline in performance once the critical information exceeds the range of global tokens and local windows.
@@
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Number Retr.KV Avg.
@@
Fig. 1b and 10 shows the latency and breakdown of MInference across different context windows on a single A100. At 100K, 300K, 500K, and 1M tokens, our method achieves speedups of 1.8$\times$, 4.1$\times$, 6.8$\times$, and 10$\times$, respectively. It reduces the pre-filling latency from 30 mins to 3 mins on a single A100 for a prompt of 1M token. By further utilizing tensor parallel [46] and context parallel [50, 31], this latency can be reduced to 40 seconds on 8x A100 GPUs. This significantly lowers the deployment cost of long-context LLMs and enhances user experience. And since our kernel is implemented based on Triton, it can be easily ported to other devices and achieve similar speedups, such as on the H100. Additionally, analyzing the latency breakdown, we found about 5%-20% of the overhead is spent on dynamic sparse index building, while the remaining time is spent on dynamic sparse calculation.
@@
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Number Retr.KV Avg.
@@
We also combined MInference with a state-of-the-art KV cache compression method SnapKV [43], as shown in Table 5. This proves our method is compatible with KV cache compression techniques. For most tasks, performance remains nearly unchanged, with the average score even showing a slight increase, which further demonstrates the potential practical value of our method as an optimization for serving long-context LLMs.
@@ 5 Related Works
Recent studies [21] have tackled the high computational cost of attention and substantial KV cache storage in long-context scenarios from two angles: pre-filling and decoding. Pre-filling optimizations are primarily categorized as State Space Models [23, 22], linear attention methods [70, 57], memory-based methods [53], hybrid methods [44, 27, 64], and prompt compression methods [41, 32, 33, 62]. However, these approaches require training from scratch or additional overhead and are difficult to implement directly in pretrained long-context LLMs. Recently, some studies [52, 83] have focused on using kNN or cluster-based sparse attention to accelerate LLM inference. However, these methods often lead to reduced accuracy, limited speedup, or are restricted to CPU scenarios.
In contrast, optimizations for the decoding stage are divided into: 1) Reusing attention KV to reduce KV cache storage [73, 3, 71, 12]; 2) Static KV cache compression patterns [82, 29]; 3) Dynamic KV cache compression patterns, including completely discarding the KV cache after compression [89, 42, 25, 56], and offloading-based methods [63, 43, 15]; 4) Cluster-based KV cache compression methods [54, 78]; 5) Methods for restoring performance loss due to KV cache compression [1, 18]; 6) Hierarchical speculative decoding methods [69]. Nevertheless, these methods do not address the heavy computational burden of the attention in the pre-filling stage.
@@ 6 Conclusion
This paper addresses the expensive computational cost and the unacceptable latency of the attention calculations in the pre-filling stage of long-context LLMs. We propose MInference, a method that accelerates the pre-filling stage by leveraging dynamic sparse attention with spatial aggregation patterns. Specifically, we categorize attention heads into three types: A-shape, Vertical-Slash, and Block-Sparse. Using a kernel-aware optimal sparse pattern search method, we identify the optimal pattern for each head. Subsequently, we utilize a fast approximation approach to build dynamic sparse masks for different inputs, and then apply these mask to perform sparse attention calculations. Experimental results on benchmarks such as InfiniteBench, RULER, language modeling, and Needle In A Haystack demonstrate that our method effectively maintains the long-context capabilities of LLMs while achieving up to a 10x speedup, reducing the latency from 30 minutes to 3 minutes per prompt for 1 million token prompts on a single A100 GPU. Additionally, we have found that similar dynamic sparse attention patterns also exist in both multi-modal LLMs [80] and encoder-decoder LLMs [66]. Using MInference for pre-filling stage inference acceleration holds great promise.
@@ D.3 Additional Ablation Study
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Number Retr.KV Avg.

### v2

@@ 1 Introduction
Large language models (LLMs) have entered the era of long-context processing, with some of them supporting context windows ranging from 128K to 10M tokens [27, 74, 54, 92, 2, 15]. These extended context windows enable LLMs to unlock a multitude of complex real-world applications, such as repository-level code understanding [8, 37, 65], long-document question-answering [10, 56], self-play reasoning [63], extreme-label in-context learning [56], and long-horizon agent tasks [87].
@@ 2.2 Attention Sparsity Exhibits Patterns
Vertical-Slash (VS) pattern The attention weights are concentrated on specific tokens (vertical lines) [59] and tokens at fixed intervals (slash lines). The positions of vertical and slash lines in this pattern dynamically change with the context content and exhibit a certain sparsity, making them difficult to be encompassed by local windows and A-shape patterns.
@@ 4 Experiments
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Num Retr.KV Avg.
@@
As shown in Table 2, MInference achieves the best overall performance on InfiniteBench compared to baseline methods. Remarkably, MInference matches or even slightly surpasses the performance of the original full attention baseline on some tasks, despite the significant acceleration it provided. From the perspective of different tasks, our method not only performs well in natural language tasks such as summarization, QA, and code, but also maintains the original model’s performance on retrieval-related tasks. Baseline methods such as StreamingLLM, on the contrary, struggle with these retrieval tasks. Additionally, on tasks such as dialogue QA, using local attention mechanisms can better handle these tasks, while our performance is closer to the original results, indicating that our method is not solely based on local windows. Extending the local windows’ intervals in StreamingLLM, i.e., w/ dilated and w/ strided, has minimal impact on the model’s performance.
@@
Comparing Fig. 1a to Fig. 6, our method effectively retains the ability to process information at different positions across various context windows, ranging from 1k to 1M tokens. In contrast, methods like StreamingLLM and InfLLM (as shown in Appendix D.1), while effective at reducing latency, experience a sharp decline in performance once critical information extends beyond the range of global tokens and local windows.
@@
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Num Retr.KV Avg.
@@
Fig. 1b and 10 shows the latency and breakdown of MInference across different context windows on a single A100. At 100K, 300K, 500K, and 1M tokens, our method achieves speedups of 1.8$\times$, 4.1$\times$, 6.8$\times$, and 10$\times$, respectively. It reduces the pre-filling latency from 30 mins to 3 mins on a single A100 for a prompt of 1M token. By further utilizing tensor parallel [50] and context parallel [55, 34], this latency can be reduced to 22 seconds on 8x A100 GPUs. This significantly lowers the deployment cost of long-context LLMs and enhances user experience. And since our kernel is implemented based on Triton, it can be easily ported to other devices and achieve similar speedups, such as on the H100 or MI300X. Additionally, analyzing the latency breakdown, we found about 5%-20% of the overhead is spent on dynamic sparse index building, while the remaining time is spent on dynamic sparse calculation.
@@
We also combined MInference with a state-of-the-art KV cache compression method SnapKV [47], as shown in Table 5. This proves our method is compatible with KV cache compression techniques. For most tasks, performance remains nearly unchanged, with the average score even showing a slight increase, which further demonstrates the potential practical value of our method as an optimization for serving long-context LLMs. This phenomenon is also observed in other works, such as ShadowKV [76].
@@
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Num Retr.KV Avg.
@@
Scaling-up on Larger LLMs
We also evaluated MInference on larger LLMs, such as LLaMA-3-70B-1M [^3] https://huggingface.co/gradientai/Llama-3-70B-Instruct-Gradient-262k. As shown in Table 6, MInference maintains strong performance even in larger models. Notably, in dynamic tasks such as KV retrieval, MInference can match or even slightly improve performance compared to full attention. In contrast, baselines like InfLLM generally struggle with tasks such as KV retrieval.
Table 6: Performance of different methods using LLaMA-3-70B-Instruct-262K on InfiniteBench [94].
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Num Retr.KV Avg.
LLaMA-3-70B-262K 20.7 10.3 84.2 9.5 14.0 33.2 61.7 97.0 100.0 34.0 46.5
StreamingLLM 20.5 8.5 52.0 10.0 12.6 27.4 61.1 14.0 10.0 0.0 21.6
InfLLM 24.1 8.1 57.0 10.0 12.9 27.4 52.3 100.0 100.0 0.0 39.2
Ours 20.6 10.1 83.4 10.0 14.1 34.1 61.9 100.0 100.0 39.0 47.3
@@ 5 Related Works
Recent studies [24] have tackled the high computational cost of attention and substantial KV cache storage in long-context scenarios from two angles: pre-filling and decoding. Pre-filling optimizations are primarily categorized as State Space Models [26, 25], linear attention methods [78, 64], memory-based methods [58, 30], hybrid methods [48, 71], and prompt compression methods [45, 36, 35, 69]. However, these approaches require training from scratch or additional overhead and are difficult to implement directly in pretrained long-context LLMs. Recently, some studies [57, 91, 41] have focused on using kNN or cluster-based sparse attention to accelerate LLM inference. However, these methods often lead to reduced accuracy, limited speedup, or are restricted to CPU scenarios.
In contrast, optimizations for the decoding stage are divided into: 1) Reusing attention KV to reduce KV cache storage [81, 3, 79, 15, 60]; 2) Static KV cache dropping [90, 32]; 3) Dynamic KV cache dropping [97, 46, 28, 62, 47, 4]; 4) Dynamic KV cache offloading [70, 18, 86, 41, 11, 76]; 5) Methods for restoring performance loss due to KV cache compression [1, 21]; 6) Hierarchical speculative decoding methods [77, 12]; 7) KV cache quantitation [53]. Nevertheless, these methods do not address the heavy computational burden of the attention in the pre-filling stage.
@@ 6 Conclusion
This paper addresses the expensive computational cost and the unacceptable latency of the attention calculations in the pre-filling stage of long-context LLMs. We propose MInference, a method that accelerates the pre-filling stage by leveraging dynamic sparse attention with spatial aggregation patterns. Specifically, we categorize attention heads into three types: A-shape, Vertical-Slash, and Block-Sparse. Using a kernel-aware optimal sparse pattern search method, we identify the optimal pattern for each head. Subsequently, we utilize a fast approximation approach to build dynamic sparse masks for different inputs, and then apply these mask to perform sparse attention calculations. Experimental results on benchmarks such as InfiniteBench, RULER, language modeling, and Needle In A Haystack demonstrate that our method effectively maintains the long-context capabilities of LLMs while achieving up to a 10×10$\times$10 × speedup, reducing the latency from 30 minutes to 3 minutes per prompt for 1 million token prompts on a single A100 GPU. Additionally, we have found that similar dynamic sparse attention patterns also exist in both multi-modal LLMs [88] and encoder-decoder LLMs [73]. Using MInference for pre-filling stage inference acceleration holds great promise.
@@ D.3 Additional Ablation Study
Methods En.Sum En.QA En.MC En.Dia Zh.QA Code.Debug Math.Find Retr.PassKey Retr.Num Retr.KV Avg.